OpenStudy (anonymous):

What is the equation of the line passing through (–3, –9) and parallel to the line y = x + 6 in standard form?

OpenStudy (nenadmatematika):

it's a line with the same slope 1, so equation is: (y+9)=1(x+3) so x-y-6=0